# Temporal-Spectral-Spatial Unified Remote Sensing Dense Prediction

### v1

@@ III-A Problem Formulation
Dense prediction tasks in remote sensing generally refer to performing structured semantic inference on input images at the pixel level. While these tasks vary in nature, their core objective remains consistent: to extract rich semantic information from remote sensing data that includes temporal, spectral, and spatial dimensions, and to output structured prediction results. This paper addresses three typical tasks—semantic segmentation, semantic change detection, and binary change detection—and formulates them as a unified tensor mapping problem. Let the input remote sensing image be represented as a four-dimensional tensor $X\in\mathbb{R}^{T_{1}\times C_{1}\times H_{1}\times W_{1}}$, where $T_{1}$ denotes the length of the time sequence, $C_{1}$ represents the number of spectral channels, and $H_{1}$ and $W_{1}$ denote the height and width of the image, respectively. The model output is represented as a tensor $Y\in\mathbb{R}^{T_{2}\times C_{2}\times H_{2}\times W_{2}}$, where $T_{2}$ represents the output time dimension, $C_{2}$ denotes the number of semantic categories, and $H_{2}$ and $W_{2}$ represent the spatial dimensions of the output image. In most cases, the input and output are spatially aligned, i.e., $H_{1}=H_{2}$ and $W_{1}=W_{2}$.
Based on this representation, the three tasks considered in this paper can be uniformly described as follows:
  • Semantic Segmentation: The model performs pixel-level multi-class classification on a remote sensing image at a given time point. This corresponds to the setup where $T_{1}=1$, $T_{2}=1$, and $C_{2}\geq 2$.
  • Semantic Change Detection: The model predicts semantic labels for each time point to analyze semantic changes between any two time points. This corresponds to $T_{2}=T_{1}$, with $C_{2}\geq 2$.
  • Binary Change Detection: The model detects changes between consecutive time points without distinguishing specific categories of change. This corresponds to $T_{2}=T_{1}-1$ and $C_{2}=2$.
It is important to note that the input data may differ significantly in terms of time length $T_{1}$, spectral channels $C_{1}$, and spatial scale $(H_{1},W_{1})$, depending on the task type or dataset. To address this variability, we propose a unified modeling framework that aims to learn a mapping function $f_{\theta}$, such that
$f_{\theta}:\mathbb{R}^{T_{1}\times C_{1}\times H_{1}\times W_{1}}\rightarrow\mathbb{R}^{T_{2}\times C_{2}\times H_{2}\times W_{2}},$ (1)
which can adapt to any input configuration and output dense prediction results that meet the requirements of the target task. This abstract modeling provides the mathematical foundation for the modular design of the subsequent methods.
III-B Overview of the Temporal-Spectral-Spatial Unified Network (TSSUN)
The proposed Temporal-Spectral-Spatial Unified Network (TSSUN) is designed as a unified modeling framework capable of addressing diverse dense prediction tasks across varying data modalities and spatial resolutions, as illustrated in Figure 2. TSSUN employs the Temporal-Spectral-Spatial Unified Strategy (TSSUS), which systematically integrates input alignment, feature extraction, and output decoding into a cohesive processing pipeline.
In the input stage, TSSUN introduces the Spectral-Spatial Unified Module (SSUM), which encodes heterogeneous spectral and spatial data into a unified representation. By leveraging spectral and spatial priors through a guided alignment mechanism, SSUM ensures data consistency across spectral and spatial dimensions, establishing a coherent input foundation for subsequent feature modeling. During the feature extraction stage, TSSUN employs the Local-Global Window Attention (LGWA) mechanism as the primary feature modeling unit, effectively capturing multi-scale features in TSS inputs. LGWA integrates three distinct window-based attention blocks, each tailored to capture local and global features at different scales. This structure balances computational efficiency and representational power, facilitating the extraction of spatial structures across multiple scales. At the encoder-decoder junction, TSSUN incorporates the Temporal Unification Module (TUM), which focuses on temporal feature fusion. TUM mitigates temporal redundancy and adjusts the temporal length of the output based on task-specific requirements, ensuring temporal coherence in the generated predictions. In the output stage, TSSUN reintroduces SSUM in a mirrored configuration to further reinforce spectral-spatial consistency during the decoding process. By progressively decoding spectral and spatial dimensions, SSUM enables unified modeling for both semantic segmentation with arbitrary class numbers and multi-resolution spatial reconstruction, thereby enhancing TSSUN’s capability to handle multi-task, multi-modal data effectively.
Overall, TSSUN leverages TSSUS to comprehensively integrate the input, feature extraction, and output stages, establishing a unified framework for heterogeneous data representation and dense prediction. The LGWA module further enhances the model’s capacity to effectively capture both local and global contextual information, ensuring a balanced representation of multi-scale and multi-modal features. The following sections provide a detailed analysis of the two core components, TSSUS and LGWA, including their structural design and functional mechanisms.
@@ III-B Overview of the Temporal-Spectral-Spatial Unified Network (TSSUN)
Figure 2: Overview of the TSSUN architecture, comprising the Spectral-Spatial Unified Module (SSUM) for input alignment, Local-Global Window Attention (LGWA) blocks for feature extraction, and the Temporal Unified Module (TUM) for temporal fusion and output adaptation. The top right panel details the Dimension Unified Module (DUM), while the bottom right panel illustrates the LGWA block structure.
III-C Temporal-Spectral-Spatial Unified Strategy (TSSUS)
The Temporal-Spectral-Spatial Unified Strategy (TSSUS) is designed to decouple and unify the representation of inputs and outputs across temporal, spectral, and spatial dimensions. In the input stage, TSSUS leverages the inherent characteristics of spectral and spatial dimensions to achieve unified encoding of remote sensing imagery. Specifically, in the spectral dimension, TSSUS captures the continuity among spectral bands and incorporates spectral metadata to ensure spectral alignment. In the spatial dimension, TSSUS extracts spatial scale representations by modeling the similarity among neighboring regions and incorporating spatial metadata. By retaining the temporal dimension at the input stage, TSSUS enables independent feature extraction for remote sensing images across different time instances. At the encoder-decoder interface, TSSUS performs feature-level fusion along the temporal axis, effectively suppressing irrelevant information while adaptively adjusting the output temporal length to accommodate dense prediction tasks. In the output stage, TSSUS further leverages the structured characteristics of dense prediction tasks in remote sensing across spectral and spatial dimensions. In the spectral dimension, it models inter-class relationships to facilitate predictions with arbitrary class numbers. In the spatial dimension, it enhances the model’s generalization capability for multi-resolution data outputs by leveraging the spatial structure consistency among neighboring regions.
@@ III-C Temporal-Spectral-Spatial Unified Strategy (TSSUS)
To facilitate the mapping from the original data space to the unified feature space across different variables (temporal, spectral, and spatial), TSSUS introduces the Dimension Unified Module (DUM), which serves as the foundational implementation for both the SSUM and the TUM. DUM employs metadata from TSS variable subsets to generate adaptive weights and biases for linear layers, thereby enabling adaptive feature mapping. As illustrated in the top-right subfigure of Figure 2, given an input tensor $X\in\mathbb{R}^{D_{1}\times H\times W}$ and its metadata $M\in\mathbb{R}^{D_{1}}$, where $D_{1}$ varies across different tasks and scenarios, DUM maps the input tensor to a unified feature $Y\in\mathbb{R}^{L\times D_{2}}$ through a hypernetwork generation process. The detailed procedure is as follows:
@@
  • Metadata Embedding: Metadata $M$ is subjected to positional encoding and tokenization. A learnable class token [CLS] is prepended to the token sequence $T\in\mathbb{R}^{(D_{1}+1)\times d}$, where $d$ denotes the embedding dimension.
@@
  • Cross-Variable Relationship Modeling: The token sequence is processed through multiple transformer blocks to capture the underlying relationships among metadata tokens.
@@
  • Adaptive Parameter Generation: The [CLS] token is linearly projected to generate bias parameters $b\in\mathbb{R}^{D_{2}}$, while the remaining tokens are projected to produce the weight matrix $W\in\mathbb{R}^{D_{1}\times D_{2}}$. The resulting $W$ and $b$ form a linear layer that maps the input features with $D_{1}$ channels to output features with $D_{2}$ channels. The input $X$ is first reshaped from $(D_{1},H,W)$ to $(L,D_{1})$, where $L=H\times W$, and then processed through the generated linear layer, resulting in an output of shape $(L,D_{2})$.
The DUM structure not only adaptively maps any TSS variable subset to a unified feature in terms of shape, but also effectively preserves the relationships between remote sensing data in original space, thereby enhancing the model’s generalization capability across heterogeneous data.
III-D Local-Global Window Attention (LGWA)
@@
Transformer networks have become a prevalent structure in remote sensing tasks due to their exceptional capacity for modeling long-range dependencies. However, the computational complexity of traditional transformers increases quadratically with sequence length, posing significant challenges for the efficient processing of large-scale remote sensing data. To address this issue, this study proposes the Local-Global Window Attention (LGWA) module, which balances computational efficiency and modeling capability by capturing fine-grained features within local windows while simultaneously incorporating global contextual information.
@@
LGWA adopts three selectable window attention shapes with varying sizes and configurations, as illustrated in Figure 3. These window shapes not only determine the feature sensitivity within each window but also directly impact computational efficiency. Therefore, selecting appropriate window configurations in different network blocks allows for an optimal trade-off between modeling effectiveness and computational resources. For a specific attention window, the input sequence $X\in\mathbb{R}^{L\times d_{M}}$ is projected into query, key, and value matrices $Q,K,V$ through linear projections, formulated as:
@@
Figure 3: Architecture of the proposed Local-Global Window Attention (LGWA)-based Transformer block, employing multiple window shapes to enhance feature extraction.
Next, attention scores are computed using the scaled dot-product attention mechanism:
@@
By adopting this modular design, the LGWA module effectively captures fine-grained features within local windows while integrating global contextual information. The flexible configuration of window shapes further optimizes computational efficiency, enabling the module to handle multi-scale remote sensing scenarios effectively.
@@ IV Experiments and Results
To validate the adaptability of TSSUN to arbitrary Temporal-Spectral-Spatial inputs and its capability to concurrently perform semantic segmentation, semantic change detection, and binary change detection tasks with support for a variable number of output classes, we conducted experiments on a total of six datasets across building and Land Cover/Land Classification (LULC) scenarios. For each scenario, a unified TSSUN model was trained using the combined training sets from all datasets within that scenario, and its performance was evaluated on their respective test sets.
In the building scenario, we selected the WHU, WHU-CD, LEVIR-CD, and TSCD datasets. While these datasets share the same number of channels in their input and label data, they exhibit variations in temporal and spatial dimensions, leading to differences in dense prediction task types. Specifically, the WHU dataset corresponds to a single-temporal building semantic segmentation task. The WHU-CD dataset is associated with a bi-temporal building semantic change detection task. The LEVIR-CD dataset is used for a bi-temporal building binary change detection task. The TSCD dataset pertains to a multi-temporal building binary change detection task, as summarized in Table I.
In the LULC scenario, we chose the LoveDA Urban and Dynamic EarthNet datasets. These datasets differ in their temporal, spectral dimensions for both input and label data. The LoveDA Urban dataset corresponds to a single-temporal semantic segmentation task, while the Dynamic EarthNet dataset is used for multi-temporal semantic change detection, semantic segmentation and binary change detection tasks, as detailed in Table I.
The experiments across multiple datasets in the building scenario were primarily designed to verify TSSUN’s adaptability to inputs and outputs with arbitrary temporal-spatial dimensions within this context. Furthermore, these experiments aimed to confirm its ability to simultaneously handle semantic segmentation, semantic change detection, and binary change detection tasks. The datasets in the LULC scenario were mainly used to validate TSSUN’s adaptability to inputs and outputs with arbitrary temporal-spectral dimensions. Additionally, these experiments were intended to demonstrate its capacity to concurrently manage semantic segmentation, binary change detection and semantic change detection tasks while supporting a flexible number of output classes.
@@
Name Scene Task Input T Output T Input C Output C Image Size Resolution Images
WHU Building SS 1 1 3 1 512×512 0.3 8189
WHU-CD Building SCD 2 2 3 1 32207×15354 0.075 1
LEVIR-CD Building BCD 2 1 3 1 1024×1024 0.5 445
TSCD Building BCD 4 3 3 1 256×256 0.5 2700
LoveDA Urban LULC SS 1 1 3 7 1024×1024 0.3 5987
DynamicEarthnet LULC SCD&SS&BCD 24 24 4 6 1024×1024 3 54750

### v2

@@ III-A Problem Formulation
Dense prediction in remote sensing encompasses a range of tasks that, despite their distinct objectives, share a common foundation: inferring structured, pixel-wise semantic information from multi-dimensional earth observation data. We formalize these tasks within a unified mathematical framework, conceptualizing them as a highly flexible tensor-to-tensor mapping problem. This abstraction is essential for accommodating the inherent heterogeneity of remote sensing data in the spatial, temporal, and spectral dimensions.
Let an input remote sensing data instance be represented by a primary data tensor $X\in\mathbb{R}^{T_{1}\times C_{1}\times H_{1}\times W_{1}}$, while the corresponding model output is a prediction tensor $Y\in\mathbb{R}^{T_{2}\times C_{2}\times H_{2}\times W_{2}}$, which is the same as the formulation in Section I. $Y$ are typically aligned with the $X$ by formulation ($H_{2}=H_{1},W_{2}=W_{1}$) and specific dense prediction task, where the spatial, temporal, and spectral dimensions of $X$ and $Y$ can vary significantly, while $T_{1}$ and $T_{2}$ is strongly correlated with the dense prediction tasks, and $C_{2}$ is strongly correlated with the predicted set of semantic categories.
To address the profound variability in data characteristics, task and category requirements, we augment this core tensor representation with explicit metadata. The model’s behavior is conditioned not only on the data tensor $X$ but also on two metadata sets: an input set $M_{in}$ describing the source data, and an output set $M_{out}$ specifying the target prediction structure. These sets are decomposed as follows:
  • Input Dimension Metadata: $M_{in}=\{M_{spa}^{in},M_{tem}^{in},M_{spe}^{in}\}$. This set provides essential context about the input data $X$. $M_{spa}^{in}$ encodes spatial information of input data, including pixel locations and spatial resolution. $M_{tem}^{in}$ contains the acquisition timestamps for each of the $T_{1}$ temporal slices. $M_{spe}^{in}$ specifies the wavelength for each of the $C_{1}$ spectral channels.
  • Output Dimension Metadata: $M_{out}=\{M_{spa}^{out},M_{tem}^{out},M_{spe}^{out}\}$. This set instructs the model on the desired output format and task. $M_{spa}^{out}$ defines the pixel locations and spatial resolution of the prediction $Y$. $M_{tem}^{out}$ determines the temporal nature of the task. It includes the target temporal length $T_{2}$ and a trainable task embedding that specifies which dense prediction task to perform. $M_{spe}^{out}$ defines the prediction’s semantic space. It consists of a dynamically selectable subset of trainable embeddings corresponding to the $C_{2}$ target classes.
Within this framework, our goal is to learn a single, unified mapping function $f_{\theta}$ parameterized by $\theta$, which can adapt to arbitrary input and output configurations. The general mapping is defined as:
$Y=f_{\theta}(X,M_{in},M_{out}).$ (1)
The three core tasks of semantic segmentation, binary change detection, and semantic change detection and various semantic category sets are thus handled as specific instances of this general function. This metadata-driven formulation provides the mathematical basis for a model that is not constrained to fixed data structures, predefined tasks or predefined category sets, enabling a truly unified approach to dense prediction in remote sensing.
III-B Overview of the Spatial-Temporal-Spectral Unified Network
@@
Figure 2: Overview of the STSUN architecture, comprising the Spectral-Spatial Unified Module for input and output alignment, the Basic blocks for feature extraction, and the Temporal Unified Module for temporal fusion and output adaptation.
The Spatial-Temporal-Spectral Unified Network unifies data representation by leveraging metadata from input and output data across spatial, temporal, and spectral dimensions. It employs trainable task embeddings and class embeddings to specify the particular dense prediction task and the semantic classes set. This design enables STSUN to adapt to various input and output shape configurations and unifies semantic segmentation, binary change detection, and semantic change detection tasks, supporting predictions across diverse class sets. STSUN primarily consists of five stages, as illustrated in Figure 2.
First, at the input stage, the spatial-spectral unified module (SSUM) leverages the input spatial metadata $M_{spa}^{in}$ and spectral metadata $M_{spe}^{in}$ for unified encoding. The temporal dimension, being intrinsically linked to the specific dense prediction task, necessitates a distinct processing approach from the spatial and spectral dimensions. If the temporal dimension were to be unified at this stage, it would be equivalent to a pixel-level fusion of multi-temporal remote sensing images. This would introduce significant interference from irrelevant information, thereby degrading the model’s performance on dense prediction tasks [13, 12]. Therefore, at this stage, the independence of the temporal dimension is preserved while the spatial and spectral dimensions are unified. In the spatial dimension, the spatial resolutions of the input images and the spatial position of each pixel are utilized as metadata for the input spatial dimension. This enables a uniform representation of data based on the similarity of spatial proximity, accommodating varied resolutions. Concurrently, in the channel dimension, the spectral wavelength of each spectral band is incorporated as metadata for the input channel dimension, facilitating a unified data representation grounded in spectral continuity.
Second, at the encoder stage, the encoder local global blocks are used to extract features from input data. As the temporal independence of remote sensing images is explicitly preserved, a shared-weight encoder then processes each temporal remote sensing image independently, extracting local and global features across distinct temporal instances within encoder local global blocks.
Third, at the encoder–decoder junction, the temporal unified module (TUM) utilizes the input temporal metadata $M_{tem}^{in}$ to fuse features from different temporal instances and adjusts these features using the output temporal metadata $M_{tem}^{out}$ to adapt them for the specific dense prediction task. The temporal position within the remote sensing image time series is used as metadata for the input time dimension, which enables feature-level fusion of remote sensing image features across various time points, effectively mitigating interference from irrelevant temporal information. Simultaneously, to guide the dense prediction task, multiple predefined trainable task embeddings are introduced. A specific task embedding is selected and incorporated as metadata with output temporal length for the output time dimension, thereby instructing the model on the required dense prediction task. It is crucial to note that operations performed in this part maintain temporal independence and do not merge with the channel dimension [13, 12]. This approach allows different dense prediction tasks to be unified within an independent temporal dimension, preventing interference from the remote sensing image features. Instead, it guides the subsequent decoder to optimize these features for the specified task, which is conducive to the model’s ability to jointly model multiple dense prediction tasks.
Fourth, the decoder subsequently processes the fused multi-temporal remote sensing image features. Since the features from multiple temporal instances have been fused and the model has been instructed on the dense prediction task via the task embedding, the shared-weight decoder optimizes the features for the specific task by removing irrelevant information, progressively aligning them with the target output.
Fifth, in the output stage, the SSUM utilizes the output spatial metadata $M_{spa}^{out}$ to restore the features to the original image dimensions and leverages the output spectral metadata $M_{spe}^{out}$ to guide the features toward making dense predictions for a specific subset of semantic categories. In the spatial dimension, the image size of the remote sensing images and the pixel position information are again leveraged as metadata for the output spatial dimension to accurately restore features to the original input image size. In the channel dimension, multiple predefined trainable category embeddings are established. From these, a subset of category embeddings is dynamically selected and integrated as metadata for the output channel dimension, explicitly indicating the set of categories the model is instructed to predict, thus enabling flexible class-agnostic prediction.
Through the arrangement of spatial, temporal, and spectral dimensions across these five stages, and the design of metadata for each dimension, STSUN accommodates diverse input/output dimensional configurations and various subsets of prediction classes. Moreover, these designs enables the effective joint modeling of semantic segmentation, binary change detection, and semantic change detection. By uncovering the correlations and complementarities among these tasks, STSUN leverages their combined data to enhance its performance across multiple dense prediction challenges.
III-C Dimension Unified Module
Figure 3: Architecture of the proposed Dimension Unified Module.
@@ III-C Dimension Unified Module
To facilitate the mapping from the original data space to the unified feature space across different dimensions (spatial, temporal and spectral), and transform the unified features into output of appropriate shapes according to the specific task and category requirements, STSUN introduces the Dimension Unified Module (DUM), which serves as the fundamental implementation of both SSUM and TUM. DUM employs metadata from STS variable subsets and trainable embeddings to generate adaptive weights and biases for linear layers, thereby enabling adaptive and metadata-aware feature transformation. As illustrated in the Figure 3, given an input tensor $X\in\mathbb{R}^{D_{1}\times H\times W}$ and its metadata $M\in\mathbb{R}^{D_{1}}$, as well as trainable embeddings $E\in\mathbb{R}^{D_{1}}$ when performing on output temporal and spectral dimension, where $D_{1}$ can refer to spatial, temporal or spectral dimension and varies across different tasks and scenarios, and $M$ is an metadata of $\{M_{spa}^{in},M_{tem}^{in},M_{spe}^{in},M_{spa}^{out},M_{tem}^{out},M_{spe}^{out}\}$. DUM maps the varying input tensor to a flexible feature $Y\in\mathbb{R}^{L\times D_{2}}$ through a hypernetwork generation process. The detailed procedure is as follows:
@@
  • Metadata Embedding: Metadata $M$ is subjected to tokenization through a linear layer and positional encoding. A learnable class token [CLS] is prepended to the token sequence $T\in\mathbb{R}^{(D_{1}+1)\times d}$, where $d$ denotes the embedding dimension.
@@
  • Cross-Variable Relationship Modeling: The metadata token sequence is processed through multiple transformer blocks to capture the underlying relationships among metadata tokens.
@@
  • Adaptive Parameter Generation: The [CLS] token is linearly projected to generate bias parameters $b\in\mathbb{R}^{D_{2}}$, while the remaining tokens are projected to produce the weight matrix $W\in\mathbb{R}^{D_{1}\times D_{2}}$. The resulting $W$ and $b$ form a linear layer that maps the input features with $D_{1}$ channels to output features with $D_{2}$ channels. The input $X$ of shape $(L,D_{1})$ can be processed through the generated linear layer, resulting in an output of shape $(L,D_{2})$. Specifically, when this module is applied to the output temporal and spectral dimensions, an additional learnable embedding $E$ is fused with $X$ prior to linear layer processing, which indicate the specific dense prediction task required for the features (corresponding to the output temporal dimension) or the specific set of semantic categories to be predicted (corresponding to the output spectral dimension).
By adaptively mapping STS variable subsets to a unified feature representation, DUM not only standardizes the feature shape but also preserves important relationships present in the original remote sensing data. Additionally, DUM can transform this unified feature representation into an output of a specific shape as required by the task, based on the dimensionality metadata of the output data. This enables STSUN to effectively support inputs and outputs with various dimensional settings, allowing it to generalize to diverse dense prediction tasks and class sets.
III-D Local-Global Window Attention
@@
Global attention mechanisms have become widely adopted in remote sensing tasks due to their strong capability to model long-range dependencies. However, their quadratic computational complexity with respect to sequence length poses a major challenge when dealing with large-scale remote sensing data. Moreover, global attention primarily captures holistic contextual features, often neglecting local spatial patterns, which are equally important for dense prediction tasks. To address this issue, this study proposes the Local-Global Window Attention module (LGWA), which improves both computational efficiency and modeling capability by jointly capturing fine-grained local features and global contextual information.
Figure 4: Architecture of the proposed LGWA-based Basic block, employing multiple local window to extract local feature and single global window to extract global feature.
@@
LGWA adopts three selectable window attention shapes with varying sizes and configurations, as illustrated in Figure 4. These configurations determine the receptive field and sensitivity of the attention operation within each window and directly influence computational cost. Selecting appropriate window configurations in different layers allows for an optimal trade-off between feature richness and resource efficiency. For a specific attention window, the input sequence $X\in\mathbb{R}^{L\times d_{M}}$ is projected into query, key, and value matrices $Q,K,V$ through linear projections, formulated as:
@@
The attention scores are computed using the scaled dot-product attention mechanism:
@@
Through this modular design, LGWA effectively captures both local structural details and global semantic context. The flexible design of window configurations allows the module to balance modeling power and computational efficiency, making it well-suited for multi-scale, high-resolution remote sensing applications.
@@ IV Experiments and Results
To validate the adaptability of STSUN to arbitrary Spatial-Temporal-Spectral inputs and outputs, and its capability to concurrently perform semantic segmentation, binary change detection, and semantic change detection tasks with support for variable class subsets, we conducted experiments on a total of six datasets across building and Land Use/Land Cover (LULC) scenarios. For each scenario, a unified STSUN model was trained using the combined training sets from all datasets within that scenario, and its performance was evaluated on their respective test sets.
In the building scenario, we selected the WHU, WHU-CD, LEVIR-CD, and TSCD datasets. While these datasets share the same number of channels in their input and label data, they exhibit variations in temporal and spatial dimensions, leading to differences in dense prediction task types, thus verifying that STSUN can adapt to various spatiotemporal input and output settings, and can unify semantic segmentation, binary change detection, and semantic change detection tasks. Specifically, the WHU dataset corresponds to a single-temporal building semantic segmentation task, the WHU-CD dataset is associated with a bi-temporal building semantic change detection task, the LEVIR-CD dataset is used for a bi-temporal building binary change detection task, and the TSCD dataset pertains to a multi-temporal building binary change detection task, as summarized in Table I.
In the LULC scenario, we chose the LoveDA Urban and Dynamic EarthNet datasets. These datasets differ in their temporal and spectral dimensions for both input and label data, which verifies that STSUN can adapt to various spatio-temporal and spectral input and output settings, unify semantic segmentation, binary change detection, and semantic change detection tasks, and support dense prediction tasks with varying semantic class sets. The LoveDA Urban dataset corresponds to a single-temporal semantic segmentation task with an semantic category subset of 7 LULC classes, while the Dynamic EarthNet dataset is used for multi-temporal semantic segmentation, binary change detection and semantic change detection tasks with another semantic category subset of 6 LULC classes, as detailed in Table I.
@@
Name Scenario Task $T_{1}$ $T_{2}$ $C_{1}$ $C_{2}$ Image Size Resolution Images
WHU [37] Building SS 1 1 3 1 512×512 0.3 8189
WHU-CD [37] Building SCD 2 2 3 1 1024×1024 0.075 480
LEVIR-CD [38] Building BCD 2 1 3 1 1024×1024 0.5 445
TSCD [39] Building BCD 4 3 3 1 256×256 0.5 2700
LoveDA Urban [40] LULC SS 1 1 3 7 1024×1024 0.3 5987
Dynamic Earthnet [41] LULC SCD &SS &BCD 24 24 4 6 1024×1024 3 54750